what is the number of molecules present in 0.062 mol of acetic acid, CH₃COOH


1
Expert's answer
2021-05-31T02:24:49-0400

1 mole contains "6.022 \times 10^{23}" atoms

Hence,

"0.062" mole contains "= 0.062 \times 6.022 \times 10^{23}" "=0.37 \times 10^{23}"